OpenStudy (anonymous):

can anyone solve this one, x^4-x^3-42x^2=0, so far i have x^2(x^2-x-42)=0

OpenStudy (turingtest):

good, so x=0 is one solution. Now just solve the quadratic using the quadratic formula

OpenStudy (anonymous):

(x^2)(x^2-x-42) = 0 (x^2)(x+6)(x-7) = 0 x = 0, 0, -6, 7 x = 0, -6, 7

OpenStudy (turingtest):

oh how nice it even factors!

OpenStudy (anonymous):

x=0,-6,7

OpenStudy (anonymous):

thanks!!